OpenStudy (anonymous):

What is the slope of the line whose equation is 3x + 6y = 9?

OpenStudy (anonymous):

@girlnotonfire

OpenStudy (anonymous):

slope of a line expressed in the form y= mx+c is m in your case your equation is 3x+6y=9 or 6y=-3x+9 or y = -3/6x+9/6 or y=-1/2 x+3/2 do you know now what the slope is ?

OpenStudy (anonymous):

:D I already know the answer. its -1/2

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Awww that's nice of you, you don't have to do that :)